P/E AND STOCK PRICE Ferrell Inc. recently reported net income of $5 million. It has 800,000 shares of common stock, which currently trades at $31 a share. Ferrell continues to expand and anticipates that 1 year from now, its net income will be $7.25 million. Over the next year, it also anticipates issuing an additional 160,000 shares of stock so that 1 year from now it will have 960,000 shares of common stock. Assuming Ferrells price/earnings ratio remains at its current level, what will be its stock price 1 year from now? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Price Earning ratio = Price / EPS

EPS (Earnings per share) = Net income for shareholders / Number of shares

Net income (now) = $ 5 million

Number of shares (now) = 800,000 shares

Current price = $ 31.

Current PE ratio = Price / EPS = 31 / (5,000,000/800,000)

Current PE ratio = 31/6.25 = 4.96 times

Future PE ratio is same as Current PE ratio = 4.96 times

Future EPS = Net income / Number of shares = $ 7.25 million / 960,000 = $ 7.5520833 Per share

PE ratio = Stock Price / EPS

Stock Price = PE ratio * EPS = 4.96 * 7.5520833 = $ 37.46 (Rounded off up to nearest cent)